Southern Ocean Iron Experiment (SOFeX) Cruise
January 5 - February 26, 2002
Skip to Log Entry from the R/V Melville
January 27, 2002: Day 23
Ship: -66 22.2 S, -171 58.2 W
So. Patch In Drifter: 66 25.2 S, 171 55.8 W
So. Patch Out Drifter: 66 36.6 S, 171 46.2 W
R/V Revelle Log Entry: The weather has been beautiful today - albeit freezing cold. Work on deck and your fingers are numb in a few minutes. The air is so dry that the evaporation super-cools wet hands.
Weve been taking advantage of the good weather to move the big Individual Bulk Containers of iron around, getting ready to prepare the next tank of iron for spreading in a couple of days. Eric Drake is removing the crane hook while Craig Hunter, Pete Strutton and David Stuebe handle the tag lines. Tammy is in the crane. Each IBC has about 1000 kg of hydrous FeSO4 in it (200 kg of pure iron).
Weve also been conducting CTD/rosette sampling all day. Weve finished a series of samples in the middle of the patch and we are now starting a sampling transect across the patch.
Steve Pierce has processed the data from last nights survey to produce the first map of the South Patch. He has added arrows that show the velocity of the patch, the biggest velocities of 25 cm/s or 0.5 miles/hour - came during the 40 knot blow.
Pete Strutton is processing the underway CO2 data from last nights survey and thinks he can already see a signal that is due to enhanced photosynthesis in the iron enriched waters. Hmm well have to wait and see if that holds up. Its great news if its correct. MELVILLE reports they have found substantial CO2 drawdown in the North Patch, as well. We expect to see them tomorrow.
Happy Birthday to my son Brad - its his 18th! Well yesterday here (26th) was his birthday, but its today there (in California).Wow, this dateline business is weird.Weve been trying to figure out when hump day is/was. We think it was yesterday, but we need a mathematician.
Well, its 5 pm so it's time for a steak and glass of wine - the highlight of the week. Bye for now.
- Ken J.
Position:58 degrees, 29 minutes South, 171 degrees, 58 minutes West
R/V Melville Log Entry: We have been operating for the last couple of days in the Northern Patch area.This is an area of about 200 square kilometers infused with iron by Ken Johnsons crew aboard Revelle some twelve days ago.We are working with on a two ship operation in order to maximize the science value of a very significant set of coupled experiments both North and South of the Polar Front Zone in the Southern Ocean. For the Northern Patch this means showing up after the Revelle has departed and continuing and also initiating for the first time, some detailed studies of biological rate processes, isotopic fractionation, bacterial production, trace metal distributions, scavenging and particle flux etc. But such a strategy has some problems. Showing up on station is kind of like walking into a kitchen halfway through the dinner preparation, finding your spouse has gone, and you have to figure out whats for dinner and locate all the pieces to put it together. Lucky for you, he left a note:"Honey Buns: Gone with kids (soccer and to pick up ice cream or something frozen), mixotrophic dinoflagellates are showing up about 6, chlorophyll is beginning to rise, having some car trouble, dogs are outside. Oh yes, the pteropods and salps are coming for dinner, see if you can make something special. See ya Tuesday. XO, Pork Chop
Critical to our ability to interpret the results of these studies is the comparison between enriched and un-enriched (ambient) conditions.This means we need to be able to locate and sample both Inside Patch areas and Outside Patch areas.We were able to get the lay of the land and after a half a day of surveying, we made a very detailed and solid station at the Inside Patch station followed by a very successful Outside Patch station.A typical station plan is appended below.After the Outside Patch station, there were a couple of things left to do before we could turn in. Some follow up work remained to be done at the Inside Station, two sediment traps needed recovery and we were still fighting problems with the ships running seawater system.A long and difficult day of surveying for the patch and buoys in 20 to 25 foot seas, high winds and some very bad rolls (35 degrees) ended in a very successful recovery of the sediment traps that the Revelle had deployed on her occupation of this northern site.In spite of the distance traveled, we were unsuccessful in relocating the core of the elusive Northern Patch and had to abandon our search for the Outside Patch sediment trap array.If the buoy starts sending positional data to the satellite again, Melville or Revelle can pick it up upon return.
Here is what we are finding (keep in mind that all data are preliminary and nothing is final):Although the Revelle reported that the chlorophyll concentrations seemed to rise a little then stop, chlorophyll concentrations in the Northern Patch have subsequently risen to over 4 times the ambient level and photosynthetic efficiency is very high everywhere in the infused area. On preliminary survey it appears that the species composition has shifted since the Revelles occupation with diatoms beginning to show up and there was a significant amount of material in the sediment traps but we cannot yet directly relate this to flux. There are significant depletions of carbon dioxide that are tightly correlated with large increases in chlorophyll.
So the findings are very interesting and the stations have been very successful (when we could do them). We are currently headed South to meet the kids and see how the game is going.
- Kenneth Coale
Typical Station Plan. Names refer to deck operations. (It never goes as planned, but we are getting to be a pretty well organized bunch, I think we will really have this down by mid February)
Update Buoy Positions and Patch Geometry, Wang
Pre Station Survey/Underway sampling (SF6, FRRF, CO2, Nuts) to pick station: Cooper, Gorbunov, Zhu, Brewster
0715 Go Flo Casts to finish bottle cleaning: Gordon, Ladizinsky, Coale
0800 CTD Cast (1), Identify water column/patch structure, natural isotopes: Palomaris, Brewster, Cooper, Jones, B. Hiscock
0845 TM Rosette Cast (1), Primary Production, Nitrogen Uptake, Si production: Wang, Herndon, Tanner, Andrews, Smith, Bidigare, Selph, Ladizinsky, Roberts
0930 CTD Cast (2), Isotopes, DOP, POP: Palomaris, Bailey, Delizo, Oliver
1015 TM Rosette Cast (2), PvsE, Size Fractionation, Nitrogen Uptake: Wang, Cochlan, Brzezinsky, Bidigare, Landry, Ladizinsky, Herndon
1115 Go Flo Casts (1), Dissolved and Particulate Trace Metals: Gordon, Coale, Ladizinsky
1115 Hand Phytoplankton Tows: Tanner
1215 CTD Cast (3), Natural Isotopic Abundance: Timothy, Delizo, Bailey, Oliver
1300 Thorium\Biomarkers, In Situ Pumps: Andrews, Bidigare, Crossin
1530 Go Flo Casts (2), Dissolved and Particulate Trace Metals: Gordon, Coale, Ladizinsky
1630 TM Rosette Cast (3), Zooplankton Grazing: Wang, Landry, Ladizinsky, Selph, Bidigare
1715 Oblique Zooplankton Tows (may be moved to evening): Sheridan, Landry
1800 Particle Sediment Trap Recovery/Redeployment: Coale, Ladizinsky, Smith, Tanner, Brewster
1930 Move to Outside station/activity/ Pump holding tanks:
1930 Map Patch for the evening with underway systems, Additional casts and activities (initiate incubations, recover/deploy sediment traps, run detailed spatial surveys in vertical sections, etc